# Exploring neural oscillations during speech perception via surrogate gradient spiking neural networks

### v1

@@ Abstract
Introduction
@@
In the field of speech processing technologies, the effectiveness of training deep artificial neural networks (ANNs) with gradient descent has led to the emergence of many successful encoder-decoder architectures for automatic speech recognition (ASR), typically trained in an end-to-end fashion over vast amounts of data [1, 2, 3, 4]. Despite recent efforts [5, 6, 7, 8] towards understanding how these ANN representations can compare with speech processing in the human brain, the cohesive integration of the fields of deep learning and neuroscience remains a challenge. Nonetheless, spiking neural networks (SNNs), a type of artificial neural network inspired by the biological neural networks in the brain, present an interesting convergence point of the two disciplines. Although slightly behind in terms of performance compared to ANNs, SNNs have recently achieved concrete progress [9] on speech command recognition tasks using the surrogate gradient method [10] which allows them to be trained via gradient descent. Further work has also shown that they can be used to define a spiking encoder inside a hybrid ANN-SNN end-to-end trainable architecture on the more challenging task of large vocabulary continuous speech recognition [11]. Their successful inclusion into contemporary deep learning ASR frameworks offers a promising path to bridge the existing gap between deep learning and neuroscience in the context of speech processing. This integration not only equips deep learning tools with the capacity to engage in speech neuroscience but also offers a scalable approach to simulate spiking neural dynamics, which supports the exploration and testing of hypotheses concerning the neural mechanisms and cognitive processes related to speech.
@@
In neuroscience, various neuroimaging techniques such as electroencephalography (EEG) can detect rhythmic and synchronised postsynaptic potentials that arise from activated neuronal assemblies. These give rise to observable neural oscillations, commonly categorised into distinct frequency bands: delta (0.5-4 Hz), theta (4-8 Hz), alpha (8-13 Hz), beta (13-30 Hz), low-gamma (30-80 Hz), and high-gamma (80-150 Hz) [12]. It is worth noting that while these frequency bands provide a useful framework, their boundaries are not rigidly defined and can vary across studies. Nevertheless, neural oscillations play a crucial role in coordinating brain activity and are implicated in cognitive processes such as attention [13, 14, 15, 16], memory [17], sensory perception [18, 19] and motor function [20, 21]. Of particular interest is the phenomenon of cross-frequency coupling (CFC) which reflects the interaction between oscillations occurring in different frequency bands [14, 22]. As reviewed by Abubaker et al. [23], many studies have demonstrated a relationship between CFC and working memory performance [24, 25]. In particular phase-amplitude coupling (PAC) between theta and gamma rhythms appears to support memory integration [26, 27, 28], preservation of sequential order [29, 30, 31] and information retrieval [32]. In contrast, alpha-gamma coupling commonly manifests itself as a sensory suppression mechanism during selective attention [33, 34], inhibiting task-irrelevant brain regions [35] and ensuring controlled access to stored knowledge [36]. Finally, beta oscillations are commonly associated with cognitive control and top-down processing [37].
@@
In the context of speech perception, numerous investigations have revealed a similar oscillatory hierarchy, where the temporal organisation of high-frequency signal amplitudes in the gamma range is orchestrated by low-frequency neural phase dynamics, specifically in the delta and theta ranges [38, 39, 40, 41, 42]. These three temporal scales – delta, theta and gamma – naturally manifest in speech and represent specific perceptual units. In particular, delta-range modulation (1-2 Hz) corresponds to perceptual groupings formed by lexical and phrasal units, encapsulating features such as the intonation contour of an utterance. Modulation within the theta-range aligns with the syllabic rate (4 Hz) around which the acoustic envelope consistently oscillates. Finally, (sub)phonemic attributes, including formant transitions that define the fine structure of speech signals, correlate with higher modulation frequencies (30-50 Hz) within the low-gamma range. The close correspondence between the perception of (sub)phonemic, syllabic and phrasal attributes on one hand, and the manifestation of gamma, theta and delta neural oscillations on the other, was notably emphasised by Giraud and Poeppel [40]. These different levels of temporal granularity inherent to speech signals therefore appear to be processed in a hierarchical fashion, with the intonation and syllabic contour encoded by earlier neurons guiding the encoding of phonemic features by later neurons. Some insights about how phoneme features end up being encoded in the temporal gyrus were given by Mesgarani et al. [43]. Drawing from recent research [44] on the neural oscillatory patterns associated with the sentence superiority effect, it is suggested that such low-frequency modulation may facilitate automatic linguistic chunking by grouping higher-order features into packets over time, thereby contributing to enhanced sentence retention. The engagement of working memory in manipulating phonological information enables the sequential retention and processing of speech sounds for coherent word and sentence representations. Additionally, alpha modulation has also been shown to play a role in improving auditory selective attention [45, 46, 47], reflecting the listener’s sensitivity to acoustic features and their ability to comprehend speech [48].
@@
Computational models [41, 49] have shown that theta oscillations can indeed parse speech into syllables and provide a reliable reference time frame to improve gamma-based decoding of continuous speech. These approaches [41, 49] implement specific models for theta and gamma neurons along with a distinction between inhibitory and excitatory neurons. The resulting networks are then optimised to detect and classify syllables with very limited numbers of trainable parameters (10-20). In contrast, this work proposes to utilise significantly larger end-to-end trainable multi-layered architectures (400k-20M trainable parameters) where all neuron parameters and synaptic connections are optimised to predict sequences of phoneme/subword probabilities, that can subsequently be decoded into words. By avoiding constraints on theta or gamma activity, the approach allows us to explore which forms of CFC naturally arise when solely optimising the decoding performance. Even though the learning mechanism is not biologically plausible, we expect that a model with sufficiently realistic neuronal dynamics and satisfying ASR performance should reveal similarities with the human brain. We divide our analysis in two parts,
@@
    Oscillations: Here we delve into the core part of our analysis on the emergence of neural oscillations within our model. Each SNN layer is treated as a distinct neuron population, from which spike trains are aggregated into a population signal similar to EEG data. Through intra- and inter-layer CFC analysis, we investigate the presence of significant delta-gamma, theta-gamma, alpha-gamma and beta-gamma PAC. We also examine the impact of SFA and recurrent connections on the synchronisation of neural activity.
@@
Results
Architectural analysis

### v2

@@ Abstract
[…]
1 Keywords:
neural oscillations, spiking neural networks, speech recognition, brain-inspired computing, deep learning, surrogate gradient, spike-frequency adaptation, neuromorphic computing
2 Introduction
@@
In the field of speech processing technologies, the effectiveness of training deep artificial neural networks (ANNs) with gradient descent has led to the emergence of many successful encoder-decoder architectures for automatic speech recognition (ASR), typically trained in an end-to-end fashion over vast amounts of data (Gulati et al., 2020; Baevski et al., 2020; Li et al., 2021; Radford et al., 2023). Despite recent efforts (Brodbeck et al., 2024; Millet et al., 2022; Millet and King, 2021; Magnuson et al., 2020) towards understanding how these ANN representations can compare with speech processing in the human brain, the cohesive integration of the fields of deep learning and neuroscience remains a challenge. Nonetheless, spiking neural networks (SNNs), a type of artificial neural network inspired by the biological neural networks in the brain, present an interesting convergence point of the two disciplines. Although slightly behind in terms of performance compared to ANNs, SNNs have recently achieved concrete progress (Hammouamri et al., 2024; Sun et al., 2023; Bittar and Garner, 2022a; Yin et al., 2021) on speech command recognition tasks using the surrogate gradient method (Neftci et al., 2019) which allows them to be trained via gradient descent. Further work has also shown that they can be used to define a spiking encoder inside a hybrid ANN-SNN end-to-end trainable architecture on the more challenging task of large vocabulary continuous speech recognition (Bittar and Garner, 2022b). Their successful inclusion into contemporary deep learning ASR frameworks offers a promising path to bridge the existing gap between deep learning and neuroscience in the context of speech processing. This integration not only equips deep learning tools with the capacity to engage in speech neuroscience but also offers a scalable approach to simulate spiking neural dynamics, which supports the exploration and testing of hypotheses concerning the neural mechanisms and cognitive processes related to speech. This investigation of complex brain functions via physiologically inspired networks aligns with the work of Pulvermüller et al. (2021); Henningsen-Schomers and Pulvermüller (2022); Pulvermüller (2023), who applied biological constraints to large-scale simulations of language learning in SNNs. We complement their approach by training on more realistic speech data, albeit at the cost of some simplifications.
@@
In neuroscience, various neuroimaging techniques such as electroencephalography (EEG) can detect rhythmic and synchronised postsynaptic potentials that arise from activated neuronal assemblies. These give rise to observable neural oscillations, commonly categorised into distinct frequency bands: delta (0.5-4 Hz), theta (4-8 Hz), alpha (8-13 Hz), beta (13-30 Hz), low-gamma (30-80 Hz), and high-gamma (80-150 Hz) (Buzsaki, 2006). It is worth noting that while these frequency bands provide a useful framework, their boundaries are not rigidly defined and can vary across studies. Nevertheless, neural oscillations play a crucial role in coordinating brain activity and are implicated in cognitive processes such as attention (Fries et al., 2001; Jensen and Colgin, 2007; Womelsdorf and Fries, 2007; Vinck et al., 2013), memory (Kucewicz et al., 2017), sensory perception (Başar et al., 2000; Senkowski et al., 2007) and motor function (MacKay, 1997; Ramos-Murguialday and Birbaumer, 2015). Of particular interest is the phenomenon of cross-frequency coupling (CFC) which reflects the interaction between oscillations occurring in different frequency bands (Jensen and Colgin, 2007; Jirsa and Müller, 2013). As reviewed in Abubaker et al. (2021), many studies have demonstrated a relationship between CFC and working memory performance (Tort et al., 2009; Axmacher et al., 2010). In particular phase-amplitude coupling (PAC) between theta and gamma rhythms appears to support memory integration (Buzsáki and Moser, 2013; Backus et al., 2016; Hummos and Nair, 2017), preservation of sequential order (Reddy et al., 2021; Colgin, 2013; Itskov et al., 2008) and information retrieval (Mizuseki et al., 2009). In contrast, alpha-gamma coupling commonly manifests itself as a sensory suppression mechanism during selective attention (Foxe and Snyder, 2011; Banerjee et al., 2011), inhibiting task-irrelevant brain regions (Jensen and Mazaheri, 2010) and ensuring controlled access to stored knowledge (Klimesch, 2012). Finally, beta oscillations are commonly associated with cognitive control and top-down processing (Engel et al., 2001).
@@
In the context of speech perception, numerous investigations have revealed a similar oscillatory hierarchy, where the temporal organisation of high-frequency signal amplitudes in the gamma range is orchestrated by low-frequency neural phase dynamics, specifically in the delta and theta ranges (Canolty et al., 2006; Ghitza, 2011; Giraud and Poeppel, 2012; Hyafil et al., 2015; Attaheri et al., 2022). These three temporal scales – delta, theta and gamma – naturally manifest in speech and represent specific perceptual units. In particular, delta-range modulation (1-2 Hz) corresponds to perceptual groupings formed by lexical and phrasal units, encapsulating features such as the intonation contour of an utterance. Modulation within the theta-range aligns with the syllabic rate (4 Hz) around which the acoustic envelope consistently oscillates. Finally, (sub)phonemic attributes, including formant transitions that define the fine structure of speech signals, correlate with higher modulation frequencies (30-50 Hz) within the low-gamma range. The close correspondence between the perception of (sub)phonemic, syllabic and phrasal attributes on one hand, and the manifestation of gamma, theta and delta neural oscillations on the other, was notably emphasised in Giraud and Poeppel (2012). These different levels of temporal granularity inherent to speech signals therefore appear to be processed in a hierarchical fashion, with the intonation and syllabic contour encoded by earlier neurons guiding the encoding of phonemic features by later neurons. Some insights about how phoneme features end up being encoded in the temporal gyrus were given in Mesgarani et al. (2014). Drawing from recent research (Bonhage et al., 2017) on the neural oscillatory patterns associated with the sentence superiority effect, it is suggested that such low-frequency modulation may facilitate automatic linguistic chunking by grouping higher-order features into packets over time, thereby contributing to enhanced sentence retention. The engagement of working memory in manipulating phonological information enables the sequential retention and processing of speech sounds for coherent word and sentence representations. Additionally, alpha modulation has also been shown to play a role in improving auditory selective attention (Strauß et al., 2014b, a; Wöstmann et al., 2017), reflecting the listener’s sensitivity to acoustic features and their ability to comprehend speech (Obleser and Weisz, 2012).
@@
Computational models (Hyafil et al., 2015; Hovsepyan et al., 2020) have shown that theta oscillations can indeed parse speech into syllables and provide a reliable reference time frame to improve gamma-based decoding of continuous speech. These approaches (Hyafil et al., 2015; Hovsepyan et al., 2020) implement specific models for theta and gamma neurons along with a distinction between inhibitory and excitatory neurons. The resulting networks are then optimised to detect and classify syllables with very limited numbers of trainable parameters (10-20). In contrast, this work proposes to utilise significantly larger end-to-end trainable multi-layered architectures (400k-20M trainable parameters) where all neuron parameters and synaptic connections are optimised to predict sequences of phoneme/subword probabilities, that can subsequently be decoded into words. By avoiding constraints on theta or gamma activity, the approach allows us to explore which forms of CFC naturally arise when solely optimising the decoding performance. Even though the learning mechanism is not biologically plausible, we expect that a model with sufficiently realistic neuronal dynamics and satisfying ASR performance should reveal similarities with the human brain. We divide our analysis in two parts,
@@
    Oscillations: We then explore the central aspect of our analysis concerning the emergence of neural oscillations within our model. Each SNN layer is treated as a distinct neuron population, from which spike trains are aggregated into a population signal similar to EEG data. Through intra- and inter-layer CFC analysis, we investigate the presence of significant delta-gamma, theta-gamma, alpha-gamma and beta-gamma PAC. We also investigate how incorporating Dale’s law, SFA and recurrent connections affect the synchronisation of neural activity.
3 Materials and methods
3.1 Spiking neuron model
Physiologically grounded neuron models such as the well known Hodgkin and Huxley model (Hodgkin and Huxley, 1952) can be reduced to just two variables (FitzHugh, 1961; Morris and Lecar, 1981). More contemporary models, such as the Izhikevich (Izhikevich, 2003) and adaptive exponential integrate-and-fire (Brette and Gerstner, 2005) models, have similarly demonstrated the capacity to accurately replicate voltage traces observed in biological neurons using just membrane potential and adaptation current as essential variables (Badel et al., 2008). With the objective of incorporating realistic neuronal dynamics into large-scale neural network simulations with gradient descent training, the linear AdLIF neuron model stands out as an adequate compromise between physiological plausibility and computational efficiency. It can be described in continuous time by the following differential equations (Gerstner and Kistler, 2002),
The neuron’s internal state is characterised by the membrane potential $u(t)$ which linearly integrates stimuli $I(t)$ and gradually decays back to a resting value $u_{\text{rest}}$ with time constant $\tau_{u}\in[3,25]$ ms. A spike is emitted when the threshold value $\vartheta$ is attained, $u(t)\geq\vartheta$, denoting the firing time $t=t^{f}$, after which the potential is decreased by a fixed amount $\vartheta-u_{r}$. In the following, we will use $u_{r}=u_{\text{rest}}$ for simplicity. The second variable $w(t)$ is coupled to the potential with strength $a$ and decay constant $\tau_{w}\in[30,350]$ ms, characterising sub-threshold adaptation. Additionally, $w(t)$ experiences an increase of $b$ after a spike is emitted, which defines spike-triggered adaptation. The differential equations can be simplified as,
$\tau_{u}\,\dot{u}(t)$ $=-u(t)-w(t)+I(t)-\tau_{u}\,\sum_{f}\delta(t-t^{f})$ (3)
$\tau_{w}\,\dot{w}(t)$ $=-w(t)+a\,u(t)+\tau_{w}\,b\,\sum_{f}\delta(t-t^{f})\,.$ (4)
and redefining neuron parameters as,
$w_{t}$ $=\beta\big{(}w_{t-1}+b\,s_{t-1}\big{)}+(1-\beta)\,a\,u_{t-1}$ (6)
$s_{t}$ $=\big{(}u_{t}\geq 1\big{)}\,.$ (7)
Stability conditions for the value of the coupling strength $a$ are derived in the Supplementary Appendix. Additionally, we constrain the values of the neuron parameters to biologically plausible ranges (Gerstner and Kistler, 2002; Augustin et al., 2013),
This AdLIF neuron model is equivalent to a Spike Response Model (SRM) with continuous time kernel functions illustrated in Fig. 1. The four neuron parameters $\tau_{u}$, $\tau_{w}$, $a$ and $b$ all characterise the shape of these two curves representing the membrane potential response to an input spike and to an emitted spike respectively. A derivation of the kernel-based SRM formulation is presented in the Supplementary Appendix.
3.1.1 Spiking layers
$I_{t}^{l}=W^{l}\,s_{t}^{l-1}+V^{l}\,s_{t-1}^{l}\,.$ (9)
Here the feedforward $W^{l}\in\mathbb{R}^{N^{l-1}\times N^{l}}$ and feedback connections $V^{l}\in\mathbb{R}^{N^{l}\times N^{l}}$ are trainable parameters. Diagonal elements of $V^{l}$ are set to zero as afterspike self inhibition is already taken into account in Eq. (5). While this choice excludes autapses, which are rare but do exist in the brain, it simplifies the model for our study. Additionally, a binary mask can be applied to matrices $W^{l}$ and $V^{l}$ to limit the number of nonzero connections. Similarly, a portion of neurons in a layer can be reduced to leaky integrate-and-fire (LIF) dynamics without any SFA by applying another binary mask to the neuron adaptation parameters $a$ and $b$. Indeed, if $a=b=0$, the adaptation current vanishes $w_{t}=0$ $\forall t\in\{1,2,\ldots,T\}$ and has no more impact on the membrane potential.
3.1.2 Surrogate gradient method
The main challenge in applying stochastic gradient descent to the derived neuronal dynamics stems from the threshold operation described in Eq. (7), which has a derivative of zero everywhere except at the threshold point, where it is undefined. To address this, a surrogate derivative can be manually specified using PyTorch (Paszke et al., 2017), which enables the application of the Back-Propagation Through Time algorithm for training the resulting SNN in a manner similar to recurrent neural network (RNN) training. In this paper, we adopt the boxcar function as our surrogate function. This choice has been proven effective in various contexts (Kaiser et al., 2020; Bittar and Garner, 2022a, b) and requires minimal computational resources as expressed by the derivative definition,
3.1.3 Spike frequency regularisation
The firing rate $f^{l}_{b,n}$ of neuron $n$ in layer $l$ when processing utterance $b$ can be calculated in Hz as,
$f_{b,n}^{l}=\frac{1}{T_{b}}\sum_{t=1}^{T}s^{l}_{b,t,n}\,,$ (11)
where $T_{b}$ is the utterance duration in seconds. We regularise the firing rates of all spiking neurons between $f_{\text{min}}=0.5$ Hz and $f_{\text{max}}=f_{\text{Nyquist}}$ using the following regularisation loss,
$\mathcal{L}_{\text{reg}}=\frac{1}{B\,L}\sum_{b=1}^{B}\sum_{l=1}^{L}\frac{1}{N^{l}}\sum_{n=1}^{N^{l}}\text{ReLU}\Big{(}f_{\text{min}}-f^{l}_{b,n}\Big{)}+\text{ReLU}\Big{(}f^{l}_{b,n}-f_{\text{max}}\Big{)}\,,$ (12)
to discourage neurons from remaining silent or from firing above the Nyquist frequency.
3.2 Overview of the auditory pathway in the brain
Sound waves are initially received by the outer ear and then transmitted as vibrations to the cochlea in the inner ear, where the basilar membrane allows for a representation of different frequencies along its length (Gundersen et al., 1978). Distinct sound frequencies induce localised membrane vibrations that activate adjacent inner hair cells. These specialised sensory cells, covering the entire basilar membrane, release neurotransmitters when activated, stimulating neighbouring auditory nerve fibers and initiating the production of action potentials. Tonotopy is maintained through the conversion of mechanical motion into electric signals as each inner hair cell, tuned to a specific frequency, only affects nearby auditory nerve fibers (Saenz and Langers, 2014). The resulting spike trains then propagate through a multi-layered neural network, ultimately reaching cortical regions associated with higher-order cognitive functions such as speech recognition. Overall, the auditory system is organised hierarchically, with each level contributing to the progressively more sophisticated processing of auditory information.
3.3 Simulated speech recognition pipeline
Our objective is to design a speech recognition architecture that, while sufficiently plausible for meaningful comparisons with neuroscience observations, remains simple and efficient to ensure compatibility with modern deep learning techniques and achieve good ASR performance. We implement the overall waveform-to-phoneme pipeline illustrated in Fig. 2 inside the Speechbrain (Ravanelli et al., 2021) framework. We provide a description of each of its components here below.
@@ 3.3 Simulated speech recognition pipeline
3.3.1 Feature extractor
Mel filterbank features are extracted from the raw waveform using 80 filters and a 25 ms window with a shift of 2 ms. This procedure down samples the 16 kHz input speech signal to a 500 Hz spectrogram representation with 80 frequency bins.
3.3.2 Auditory CNN
3.3.3 Auditory nerve fibers
3.3.4 Multi-layered SNN
3.3.5 Spikes to probabilities
The spike trains of the last layer are sent to a an average pooling module which down samples their time dimension to 25 Hz. These are then projected to 512 phoneme features using two fully connected (FC) layers with Leaky-ReLU activation. A third FC layer with Log-Softmax activation finally projects them to 40 log-probabilities representing 39 phoneme classes and a blank token as required by connectionist temporal classification (CTC).
3.3.6 Training and inference
The log-probabilities are sent to a CTC loss (Graves et al., 2006) so that the parameters of the complete architecture can be updated through back propagation. Additionally, regularisation of the firing rate as defined in Eq. (12) is used to prevent neurons from being silent or firing above the Nyquist frequency. At inference, CTC decoding is used to output the most likely phoneme sequence from the predicted log-probabilities, and the phoneme error rate (PER) is computed to evaluate the model’s performance.
3.4 Physiological plausibility and limitations
3.4.1 Cochlea and inner hair cells
While some of the complex biological processes involved in converting mechanical vibrations to electric neuron stimuli can be abstracted, we assume that the key feature to retain is the tonotopic encoding of sound information. A commonly used metric in neuroscience is the ratio of characteristic frequency to bandwidth, which defines how sharply tuned a neuron is around the frequency it is most responsive to. As detailed in Supplementary Table S6, from measured Q10 values in normal hearing humans reported in Devi et al. (2022), we evaluate that a single auditory nerve fiber should receive inputs from 5-7 adjacent frequency bins when using 80 Mel filterbank features. The adoption of a Mel filterbank frontend can be justified by its widespread utilisation within deep learning ASR frameworks. Although we do not attempt to directly model cochlear and hair cell processing, we can provide a rough analogue in the form of Mel features passing through a trainable convolution module that yields plausible ranges of frequency sensitivity for our auditory nerve fibers.
3.4.2 Simulation time step
Modern ASR systems (Gulati et al., 2020; Radford et al., 2023) typically use a frame period of $\Delta t=10$ ms during feature extraction, which is then often sub-sampled to 40 ms using a CNN before entering the encoder-decoder architecture. In the brain, typical minimal inter-spike distances imposed by a neuron’s absolute refractory period can vary from 0 to 5 ms (Gerstner and Kistler, 2002). We therefore assume that using a time step greater than 5 ms could result in dynamics that are less representative of biological phenomena. Although using a time step $\Delta t<1$ ms may yield biologically more realistic simulations, we opt for time steps ranging from 1 to 5 ms to ensure computational efficiency. After the SNN, the spike trains of the last layer are down-sampled to 25 Hz via average pooling on the time dimension. This prevents an excessive number of time steps from entering the CTC loss, which could potentially hinder its decoding efficacy. We use $\Delta t=5$ ms for most of the hyperparameter tuning to reduce training time, but favour $\Delta t=2$ ms for the oscillation analysis so that the full gamma range of interest (30-150 Hz) remains below the Nyquist frequency at 250 Hz.
3.4.3 Neuron model
The LIF neuron model is an effective choice for modelling auditory nerve fibers as it accurately represents their primary function of encoding sensory inputs into spike trains. We avoid using SFA and recurrent connections, as they are not prevalent characteristics of nerve fibers. On the other hand, for the multi-layered SNN, the linear AdLIF neuron model with layer-wise recurrent connections stands out as a good compromise between accurately reproducing biological firing patterns and remaining computationally efficient (Bittar and Garner, 2022a; Deckers et al., 2024). Although less popular than the moving threshold formulation by Bellec et al. (2018), recently reviewed in Ganguly et al. (2024), our implementation of SFA using the AdLIF model combines spike-triggered adaptation with subthreshold coupling. Previous work (Bittar and Garner, 2022a) demonstrated that the AdLIF outperforms moving threshold implementations (Yin et al., 2021; Salaj et al., 2021; Shaban et al., 2021; Yin et al., 2020) in speech command recognition tasks. Nevertheless, we will still implement and train an additional model with moving threshold SFA to ensure that our conclusions hold consistently across different SFA models.
3.4.4 Organisation in layers
Similarly to ANNs, our simulation incorporates a layered organisation, which facilitates the progressive extraction and representation of features from low-order to higher-order, without the need of concretely defining and distinguishing neuron populations. This fundamental architectural principle aligns with the general hierarchical processing observed in biological brains. However, it oversimplifies the complexities of auditory processing, which extends beyond a straightforward sequential framework. While there is some sort of sequential processing in sub-cortical structures, the levels of processed features are more intricate than a simple hierarchy. This simplification is made to ensure compatibility with deep learning frameworks.
3.4.5 Layer-wise recurrence
3.4.6 Excitatory and inhibitory
In the neuroscience field, neurons are commonly categorised into two types: excitatory neurons, which stimulate action potentials in postsynaptic neurons, and inhibitory neurons, which reduce the likelihood of spike production in postsynaptic neurons. This principle is referred to as Dale’s law.
In ANNs, weight matrices are commonly initialised with zero mean and a symmetric distribution, so that the initial number of excitatory and inhibitory connections is balanced. During training, synaptic connections are updated across all layers without enforcing a distinction between excitatory and inhibitory neurons. Dale’s law can nevertheless be imposed (Li et al., 2024; Cornford et al., 2021) even if it typically results in slightly reduced performance.
In our baseline model, Dale’s law is not applied, so that similarly to standard ANNs, weight matrices are trained without constraining values to be positive or negative. Additionally, we train separate SNNs with Dale’s law to evaluate its impact on neural oscillations. In this setup, half the neurons are excitatory and half are inhibitory, while the auditory nerve fibers are all excitatory.
3.4.7 Delays
In biological neural networks, the propagation time of spikes between neurons introduces delays, primarily due to axonal transmission. To incorporate and assess the impact of these delays on neural oscillations, we additionally train separate SNNs using dilated convolutions in the temporal dimension instead of fully connected feedforward matrices. This approach, introduced in Hammouamri et al. (2024), allows us to introduce controlled delays directly into the network architecture. Based on their configuration for speech command recognition tasks, we use a maximum delay value of 300 ms.
3.4.8 Learning rule
Stochastic gradient descent, though biologically implausible due to its global and offline learning framework, allows us to leverage parallelisable and fast computations to optimise larger-scale neural networks. While this approach facilitates effective training and scaling, it diverges from biologically inspired synaptic plasticity mechanisms, such as those mediated by AMPA and NMDA receptors.
3.4.9 Decoding into phoneme sequences
Although lower PERs could be achieved with a more sophisticated decoder, our primary focus is on analysing the spiking layers within the encoder. For simplicity, we therefore opt for straightforward CTC decoding, which more directly reflects the encoder’s capabilities.
3.4.10 Hybrid ANN-SNN balance
3.5 Speech processing tasks
The following datasets are used in our study.
  • The TIMIT dataset (Garofolo et al., 1993) provides a comprehensive and widely utilised collection of phonetically balanced American English speech recordings from 630 speakers with detailed phonetic transcriptions and word alignments. It represents a standardised benchmark for evaluating ASR model performance. The training, validation and test sets contain 3696, 400 and 192 sentences respectively. Utterance durations vary between 0.9 to 7.8 seconds. Due to its compact size of approximately five hours of speech data, the TIMIT dataset is well-suited for investigating suitable model architectures and tuning hyperparameters. It is however considered small for ASR hence the use of Librispeech presented below.
  • The Librispeech corpus (Panayotov et al., 2015) contains about 1,000 hours of English speech audiobook data read by over 2,400 speakers with utterance durations between 0.8 and 35 seconds. Given its significantly larger size, we only train a few models selected on TIMIT to confirm that our analysis holds when scaling to more data.
  • The Google Speech Commands dataset (Warden, 2018) contains one-second audio recordings of 35 spoken commands such as ”yes,” ”no,” ”stop,” ”go,” ”left,” ”right” ”up”. The training, validation and testing splits contain approximately 85k, 10k and 5k examples respectively. It is used to test whether similar CFCs arise when simply recognising single words instead of phoneme or subword sequences.
[…]
3.6 Analysis methods
3.6.1 Hyperparameter tuning
  • [noitemsep]
  • number of Mel bins: 80
  • Mel window size: 25 ms
  • auditory CNN kernel size (7, 7)
  • auditory CNN stride: (1, 1)
  • auditory CNN padding: (7, 0)
  • average pooling size: 40 / $\Delta t$
  • number of phoneme FC layers: 2
  • number of phoneme features: 512
  • dropout: 0.15
  • activation: LeakyReLU
  • [noitemsep]
  • filter bank hop size controlling the SNN time step in ms: {1, 2, 5}
  • number of auditory CNN channels (filters): {8, 16, 32, 64, 128}
  • number of SNN layers: {1, 2, 3, 4, 5, 6, 7}
  • neurons per SNN layer: {64, 128, 256, 512, 768, 1024, 1536, 2048}
  • proportion of neurons with SFA: [0, 1]
  • feedforward connectivity: [0, 1]
  • recurrent connectivity: [0, 1]
3.6.2 Population signal
In the neuroscience field, EEG stands out as a widely employed and versatile method for studying brain activity. By placing electrodes on the scalp, this non-invasive technique measures the aggregate electrical activity resulting from the synchronised firing of neurons within a specific brain region. An EEG signal therefore reflects the summation of postsynaptic potentials from a large number of neurons operating in synchrony. The typical sampling rate for EEG data is commonly in the range of 250 to 1000 Hz which matches our desired simulation time steps. With our SNN, we do not have EEG signals but directly the individual spike trains of all neurons in the architecture. In order to perform similar population-level analyses, we sum the binary spike trains $s^{l}_{b}\in\{0,1\}^{T\times N^{l}}$ emitted by all neurons in a specific layer $l$ for a single utterance $b$ as follows,
$p^{l}_{b,t}=\sum_{n=1}^{N^{l}}s^{l}_{b,t,n}\,.$ (13)
Before performing the PAC analysis, the resulting population activity signal $p^{l}_{b,t}$ is then normalised over the time dimension with a mean of 0 and a standard deviation of 1, yielding the normalised population signal $\hat{p}^{l}_{b,t}$ defined as,
$\hat{p}^{l}_{b,t}=\frac{p^{l}_{b,t}-\mu^{l}_{b}}{\sigma^{l}_{b}}\,,$ (14)
where $\mu^{l}_{b}$ is the mean and $\sigma^{l}_{b}$ is the standard deviation of $p^{l}_{b,t}$ over the time dimension.
3.6.3 Phase-amplitude coupling
Using finite impulse response band-pass filters, the obtained population signals are decomposed into different frequency ranges. We study CFC in the form of PAC both within a single population and across layers. This technique assesses whether a relationship exists between the phase of a low frequency signal and the envelope (amplitude) of a high frequency signal.
As recommended in Hülsemann et al. (2019), we implement both the modulation index (Tort et al., 2008) and mean vector length (Canolty et al., 2006) metrics to quantify the observed amount of PAC. For each measure type, the observed coupling value is compared to a distribution of 10,000 surrogates to assess the significance. Surrogate couplings are computed by disrupting the temporal order of the amplitude time series while preserving its overall characteristics. Specifically, the amplitude time series is permuted by cutting it at a random point and reversing the order of the two segments. This method, as discussed by Hülsemann et al. (2019), maintains all inherent properties of the original data except for the temporal relationship between phase angle and amplitude magnitude, providing a conservative test of significance. A p-value is then obtained by fitting a Gaussian function on the distribution of surrogate coupling values and calculating the area under the curve for values greater than the observed coupling value.
As pointed out in Jones (2016), it is important to note that observed oscillations can exhibit complexities such as non-sinusoidal features and brief transient events on single trials. Such nuances become aggregated when averaging signals, leading to the widely observed continuous rhythms. We therefore perform all analysis on single utterances.
For intra-layer interactions, a single population signal is used to extract both the low-frequency oscillation phase and the high-frequency oscillation amplitude. In a three-layered architecture, these interactions include nerve-nerve, first layer-first layer, second layer-second layer, and third layer-third layer couplings.
For inter-layer interactions, we consider couplings between the low-frequency oscillation phase in one layer and the high-frequency oscillation amplitude in all subsequent layers. These interactions include nerve-first layer, nerve-second layer, nerve-third layer, first layer-second layer, first layer-third layer, second layer-third layer couplings.
For all aforementioned intra- and inter-layer combinations, we use delta (0.5-4 Hz), theta (4-8 Hz), alpha (8-13 Hz) and beta (13-30 Hz) ranges as low-frequency modulating bands, and low-gamma (30-80 Hz) and high-gamma (80-150 Hz) ranges as high-frequency modulated bands. For a given model, we iterate through the 64 longest utterances in the TIMIT test set. For each utterance, we consider the 10 aforementioned intra- and inter-layer relations, as well as the 8 possible combinations of low-frequency to high-frequency bands. We conduct PAC testing on each of the 5,120 resulting coupling scenarios, and only consider a coupling to be significant when both modulation index and mean vector length metrics yield p-values below 0.05.
4 Results
4.1 Architectural analysis